OpenStudy (anonymous):

how do you solve the square root of 78?

OpenStudy (tkhunny):

There is nothing to "solve"... One must have an equation for that. Simplfy? Look for perfect squares. Prime Factorization will do it! \(\sqrt{78} = \sqrt{2\cdot 3\cdot 13}\) -- Hmm.... Looks like we're stuck with 78.